OpenStudy (anonymous):

a nuclear cooling tower cross section is a hyperbola having a diameter of 60ft at the center. the distance between the two foci is 102ft . What is the equation for the hyperbola ?

OpenStudy (anonymous):

\[\frac{x^2}{900}-\frac{y^2}{1701}=1 \] About 60% sure on the one. Had to bail to do some chores. The attachment looks like a cooling tower.